The difference between two numbers is 14 and half of their sum is 25. What are the numbers?
Answer by solver91311(24713)

x is one number
y is the other number
Their difference is 14: x - y = 14.

Half of their sum is 25, so the whole of their sum is 50: x + y = 50

Add the two equations: (x - y)+(x + y) = 14 + 50

2x + 0y = 64

x = 32

32 - y = 14

y = 32 - 14 = 18

Check:
32 - 18 = 14

32 + 18 = 50, 50/2 = 25 Answer checks.